the sum of three consecutive multiples of 5 is 90. find the numbers.

Answer by Charles1947(18)

the sum of three consecutive multiples of 5 is 90. find the numbers.
Let the first number be x.
The next number is x+5.
The next number is x+10.
x+(x+5)+(x+10)=90
x+x+5+x+10=90
combine terms
3x+15+90
Subtract 15 from both sides of the equation.
3x=75
Divide both sides of the equation by 3
x=25, x+5=30, x+10=35
Check 25+30+35=90